# On the Foundation of Distributionally Robust Reinforcement Learning

### v3

@@ A.1 Proof of Theorem 1
First, by Lemma 1, $v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ is the smallest amongst all the 9 values in Lemma 1. This and Proposition 2 and 4 implies that
$\mu[u^{*}]=v^{*}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi_{\mathrm{M}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq\mu[u^{*}].$
Therefore, all the inequalities above are equality. Moreover, by Lemma 1 and Proposition 3,
$\mu[u^{*}]\geq v^{*}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\geq v^{*}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\geq v^{*}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})=\mu[u^{*}].$
So, equalites hold for above inequalities as well. Finally, by the above equalities and Proposition 3,
$\mu[u^{*}]=v^{*}(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi_{\mathrm{M}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq\mu[u^{*}].$
This completes the proof. ∎
A.2 Proof of Theorem 2
Now we go back to the general case where $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ is an arbitrary subset. We highlight a consequence of the extra assumption (3.2) on $u^{*}$. This strengthens the statement of Proposition 2 to the case of $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$.
@@ A.2 Proof of Theorem 2
Assume that the solution $u^{*}$ to (3.1) further satisfies (3.2), then $v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\leq\mu[u^{*}]$.
The proof to this proposition is deferred to Appendix B.6. Assume the validity of this proposition, we present our proof of item 1 in Theorem 2.
Proof of Item 1 in Theorem 2.
Note that, by Lemma 1,
$v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi,\mathrm{K})\leq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$
for all $\mu\in\mathcal{P}(\mathcal{S})$ and any mixture of $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\Pi=\Pi^{\mathrm{C}}_{\mathrm{H}},\Pi^{\mathrm{C}}_{\mathrm{M}},\Pi^{\mathrm{C}}_{\mathrm{S}}$. This, Theorem 1, and Proposition 5 implies that
$\mu[u^{*}]=v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi,\mathrm{K})\leq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\leq v^{*}(\mu,\Pi^{\mathrm{C}}_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}).$
@@
Next we prove item 2 of Theorem 2. We claim that it is sufficient to prove the following proposition:
@@
$\sup_{\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}}v(\mu,\pi,\kappa)\leq\mu[u^{*}]+\epsilon.$
We prove the sufficiency of Proposition 6 and defer its proof to the appendix Section B.6 as well.
Proof of Item 2 in Theorem 2.
Note that by assumption, elementary properties of sup-inf, and set inclusions, it is clear that
$\mu[u^{*}]=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa)\leq\inf_{\kappa\in\mathrm{K}}\sup_{\pi\in\Pi}v(\mu,\pi,\kappa)\leq\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}}\sup_{\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}}v(\mu,\pi,\kappa)$ (A.1)
for any mixture of $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\Pi=\Pi^{\mathrm{C}}_{\mathrm{H}},\Pi^{\mathrm{C}}_{\mathrm{M}},\Pi^{\mathrm{C}}_{\mathrm{S}}$. On the other hand, the statement of Proposition 6 implies that
$\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}}\sup_{\pi\in\Pi^{\mathrm{C}}_{\mathrm{H}}}v(\mu,\pi,\kappa)\leq\mu[u^{*}].$
Therefore, we conclude that the inequalities in (A.1) must be equalities. This proves item 2 of Theorem 2. ∎
A.3 Proof of Theorem 3
@@
Notice that since $\mathcal{P}_{s}=\prod_{a\in A}\mathcal{P}_{s,a}$, the Bellman equation for the $q$-function (3.4) can be written as
$q(s,a)=r(s,a)+\gamma\inf_{p\in\mathcal{P}_{s,a}}p\left[\sup_{b\in A}q(\cdot,b)\right].$ (A.2)
Since $u^{*}$ solves (3.1), we naturally have that
$u^{*}(s)$ $\leq\sup_{d\in\mathcal{P}(\mathcal{A})}\inf_{p_{s}\in\mathcal{P}_{s}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[u^{*}]$
$=\sup_{d\in\mathcal{P}(\mathcal{A})}\inf_{p_{s}\in\mathcal{P}_{s}}\sum_{a\in A}d[a]r(s,a)+\gamma d[a]p_{s,a}[u^{*}]$
=$\stackrel{{\scriptstyle(i)}}{{=}}\sup_{d\in\mathcal{P}(\mathcal{A})}\inf_{p_{s,a_{1}}\in\mathcal{P}_{s,a_{1}}}\dots\inf_{p_{s,a_{|A|}}\in\mathcal{P}_{s,a_{|A|}}}\sum_{a\in A}d[a]r(s,a)+\gamma d[a]p_{s,a}[u^{*}]$
$=\sup_{d\in\mathcal{P}(\mathcal{A})}\sum_{a\in A}d[a]r(s,a)+\gamma d[a]\inf_{p_{s,a}\in\mathcal{P}_{s,a}}p_{s,a}[u^{*}]$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\max_{a\in A}r(s,a)+\gamma\inf_{p_{s,a}\in\mathcal{P}_{s,a}}p_{s,a}[u^{*}]$
$=\max_{a\in A}q^{*}(s,a)$
On the other hand, since $\mathcal{Q}^{\mathrm{D}}\subset\mathcal{Q}$,
$u^{*}(s)$ $\geq\max_{a\in A}\inf_{p_{s}\in\mathcal{P}_{s}}r(s,a)+\gamma\delta_{\left\{{a}\right\}}\otimes p_{s}[u^{*}]$
$=\max_{a\in A}r(s,a)+\gamma\inf_{p_{s}\in\mathcal{P}_{s}}p_{s,a}[u^{*}]$
=$\stackrel{{\scriptstyle(i)}}{{=}}\max_{a\in A}r(s,a)+\gamma\inf_{p_{s,a}\in\mathcal{P}_{s,a}}p_{s,a}[u^{*}]$
$=\max_{a\in A}q^{*}(s,a)$
where $(i)$ again follows from SA-rectangularity. Therefore, $u^{*}(s)=\max_{a\in A}q^{*}(s,a)$ and hence $q^{*}$ solves (A.2).
Finally, we show that $u^{*}$ solves (3.2). By the definition of $q^{*}$, for any $\delta>0$, there exists $\psi(s,a)\in\mathcal{P}_{s,a}$ s.t.
$r(s,a)+\gamma\psi(s,a)[u^{*}]-\delta\leq q^{*}(s,a)$
for all $(s,a)\in S\times A$. Then, we consider
$u^{*}(s)$ $\leq\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{d\in\mathcal{Q}}d[r(s,\cdot)]+\gamma d\otimes p_{s}[u^{*}]$
$\leq\sup_{d\in\mathcal{P}(\mathcal{A})}d[r(s,\cdot)]+\gamma d\otimes\psi(s,\cdot)[u^{*}]$
=$\stackrel{{\scriptstyle(i)}}{{=}}\max_{a\in A}r(s,a)+\gamma\psi(s,a)[u^{*}]$
$\leq\max_{a\in A}q^{*}(s,a)+\delta$
$=u^{*}(s)+\delta$
Appendix B Proofs of Key Propositions in Sections 3 and A
A technique that we will be using in the subsequent proof is to truncate the infinite horizon reward to a finite time $T$ that is large enough. Specifically, consider for any $\pi\in\Pi_{\mathrm{H}}$ and $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$,
$v(\mu,\pi,\kappa)$ $=E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})+\gamma^{T}u^{*}(X_{T})+\sum_{k=T}^{\infty}\gamma^{k}r(X_{k},A_{k})-\gamma^{T}u^{*}(X_{T})\right]$ (B.1)
$\leq E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})+\gamma^{T}u^{*}(X_{T})\right]+\frac{\epsilon}{2}$
if $\gamma^{T}\leq\epsilon(1-\gamma)/4$. In light of (B.1), we choose $T$ so that for all $\pi\in\Pi_{\mathrm{H}}$ and $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$,
$v_{\mu}^{\pi}(T,\kappa):=E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})+\gamma^{T}u^{*}(X_{T})\right]$ (B.2)
satisfies $|v_{\mu}^{\pi}(T,\kappa)-v(\mu,\pi,\kappa)|\leq\epsilon/2$.
B.1 Proof of Lemma 1

### v4

@@ B.1 Proof of Theorem 1
Proof of Theorem 1 First, by Lemma 1, $v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ is the smallest amongst all the 9 values in Lemma 1. This and Propositions 6 and 3 imply that
$E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v(\mu,\Pi_{\mathrm{M}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq E_{\mu}[u^{*}(X_{0})].$
Therefore, all the above inequality signs achieve equalities. In addition, by Lemma 1 and Proposition 2 and 3,
$E_{\mu}[u^{*}(X_{0})]\geq v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\geq v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\geq v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})=E_{\mu}[u^{*}(X_{0})].$ (B.1)
So, all the above inequality signs must achieve equalities as well. Finally, combining (B.1) and Proposition 2,
$E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq v(\mu,\Pi_{\mathrm{M}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}})\leq E_{\mu}[u^{*}(X_{0})].$
Again, all the above inequality signs must achieve equalities. This completes the proof. ∎
B.1.1 Proof of Proposition 2
Proof.
Proof It is equivalent to show that for any $\pi\in\Pi_{\mathrm{U}}$
$\inf_{\kappa\in\mathrm{K}^{\mathrm{S}}_{\mathrm{U}}}v(\mu,\pi,\kappa)\leq E_{\mu}[u^{*}(X_{0})];$
for $\mathrm{U}=\mathrm{H},\mathrm{M},\mathrm{S}$, respectively.
Fix any $\epsilon>0$. Given a policy $\pi=(\pi_{0},\pi_{1},\dots)\in\Pi_{\mathrm{H}}$ (or $\Pi_{\mathrm{M}}$, $\Pi_{\mathrm{S}}$), we prove this by constructing an adversarial policy $\kappa=(\kappa_{0},\kappa_{1},\dots)\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ (or $\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$, $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ resp.), so that $v(\mu,\pi,\kappa)\leq E_{\mu}[u^{*}(X_{0})]+\epsilon.$
For all $k\geq 0$ and $h_{k}=(g_{k-1},s)\in\mathbf{H}_{k}$, as $u^{*}$ solves (3.1),
$\inf_{p_{s}\in\mathcal{P}_{s}}E_{\pi_{k}(\cdot|h_{k}),p_{s}}[r(s,A_{0})+\gamma u^{*}(X_{1})]\leq u^{*}(s).$
Hence, for any $\delta>0$, there always exists $\kappa_{k}(\cdot|h_{k},\cdot)\in\mathcal{P}_{s}$ so that
$\sum_{a\in A}\sum_{s^{\prime}\in S}\pi_{k}(a|h_{k})\kappa_{k}(s^{\prime}|h_{k},a)[r(s,a)+\gamma u^{*}(s^{\prime})]\leq u^{*}(s)+\delta$ (B.2)
for all $k\geq 0$ and $h_{k}=(g_{k-1},s)\in\mathbf{H}_{k}$. Moreover, by the same argument, if $\pi_{k}(\cdot|h_{k})=\pi_{k}(\cdot|s)$ is Markov (or $\pi_{k}(\cdot|h_{k})=\pi(\cdot|s)$ is Markov time-homogeneous), there exists $\kappa_{k}(\cdot|h_{k},\cdot)=\kappa_{k}(\cdot|s,\cdot)$ (or $\kappa_{k}(\cdot|h_{k},\cdot)=\kappa(\cdot|s,\cdot)$ Markov time-homogeneous) so that (B.2) holds for all $k\geq 0$ and $h_{k}=(g_{k-1},s)\in\mathbf{H}_{k}$. We construct an adversarial policy $\kappa:=(\kappa_{0},\kappa_{1},\dots)\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ (or $\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$, $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ resp.) with $\left\{{\kappa_{k}:k\geq 0}\right\}$ specified above.
With this construction, we consider
$E_{\mu}^{\pi,\kappa}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})\right]$ $=E_{\mu}^{\pi,\kappa}E_{\mu}^{\pi,\kappa}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})\middle|H_{k-1}\right]$ (B.3)
$\stackrel{{\scriptstyle(i)}}{{=}}E_{\mu}^{\pi,\kappa}\sum_{a\in A}\sum_{s^{\prime}\in S}\pi_{k}(a|H_{k-1})\kappa_{k}(s^{\prime}|H_{k-1},a)[r(X_{k-1},a)+\gamma u^{*}(s^{\prime})]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}E_{\mu}^{\pi,\kappa}u^{*}(X_{k-1})+\delta$
where $(i)$ follows from that given $H_{k-1},A_{k-1}=a$, $X_{k}\sim\kappa_{k}(\cdot|H_{k-1},a)$; $(ii)$ follows from (B.2).
We use this inequality to bound $v_{\mu}^{\pi}(T,\kappa)$ defined in (A.2). We choose $T$ large so that for all $\pi,\kappa$, $|v_{\mu}^{\pi}(T,\kappa)-v(\mu,\pi,\kappa)|\leq\epsilon/2$. Then,
$\stackrel{{\scriptstyle(i)}}{{\leq}}E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})\right]+\gamma^{T-1}\left[E_{\mu}^{\pi,\kappa}u^{*}(X_{T-1})-E_{\mu}^{\pi,\kappa}r(X_{T-1},A_{T-1})+\delta\right]$
$=v_{\mu}^{\pi}(T-1,\kappa)+\gamma^{T-1}\delta$
$\leq\dots$
$\leq v_{\mu}^{\pi}(0,\kappa)+\sum_{k=0}^{T-1}\gamma^{k}\delta$
$\leq E_{\mu}[u^{*}(X_{0})]+\frac{\delta}{1-\gamma}$
where $(i)$ uses (B.3). Since $\delta>0$ is arbitrary, we let $\delta=\epsilon(1-\gamma)/2$ and conclude that
$v(\mu,\pi,\kappa)\leq v^{\pi}_{\mu}(T,\kappa^{\prime})+\frac{\epsilon}{2}\leq E_{\mu}[u^{*}(X_{0})]+\epsilon.$
Recall that $\epsilon>0$ and $\pi\in\Pi_{\mathrm{H}}$ (or $\Pi_{\mathrm{M}}$, $\Pi_{\mathrm{S}}$) are arbitrary, with corresponding $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ (or $\mathrm{K}_{\mathrm{M}}^{\mathrm{S}}$, $\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ resp.) This completes the proof of the proposition. ∎
B.1.2 Proof of Proposition 3
Proof.
Proof
Since $u^{*}$ is the solution to (3.1), for any fixed $\delta>0$, there exists Markov time-homogeneous decision rule $\Delta:S\rightarrow\mathcal{Q}$ s.t. for all $s\in S$,
$u^{*}(s)\leq\inf_{p_{s}\in\mathcal{P}_{s}}E_{\Delta(\cdot|s),p_{s}}[r(s,A_{0})+\gamma u^{*}(X_{1})]+\delta.$ (B.4)
Let $\pi=(\Delta,\Delta,\dots)\in\Pi_{\mathrm{S}}$. We consider for any $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$,
$\gamma E_{\mu}^{\pi,\kappa}\left[u^{*}(X_{k})\right]$ $=\gamma E_{\mu}^{\pi,\kappa}E_{\mu}^{\pi,\kappa}\left[u^{*}(X_{k})\middle|G_{k-1}\right]$ (B.5)
$\stackrel{{\scriptstyle(i)}}{{=}}\gamma E_{\mu}^{\pi,\kappa}\sum_{s^{\prime}\in S}\kappa_{k-1}(s^{\prime}|G_{k-1})u^{*}(s^{\prime})$
$=\gamma E_{\mu}^{\pi,\kappa}E_{\mu}^{\pi,\kappa}\left[\sum_{s^{\prime}\in S}\kappa_{k-1}(s^{\prime}|G_{k-1})u^{*}(s^{\prime})\middle|H_{k-1}\right]$
$=\gamma E_{\mu}^{\pi,\kappa}\sum_{a\in A}\sum_{s^{\prime}\in S}\Delta(a|X_{k-1})\kappa_{k-1}(s^{\prime}|H_{k-1},a)u^{*}(s^{\prime})$
where $(i)$ is because under $E_{\mu}^{\pi,\kappa}[\cdot|G_{k-1}]$, $X_{k}\sim\kappa_{k-1}(\cdot|G_{k-1})$ and the last equality follows from that under $E_{\mu}^{\pi,\kappa}[\cdot|H_{k-1}]$, $A_{k-1}\sim\Delta(\cdot|X_{k-1})$.
Now, we define the random adversarial decision $K_{k-1}(\cdot|X_{k-1},a)=\kappa_{k-1}(\cdot|H_{k-1},a)$. By (B.5)
$E_{\mu}^{\pi,\kappa}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})\right]$ $=E_{\mu}^{\pi,\kappa}E_{\mu}^{\pi,\kappa}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})|H_{k-1}\right]$ (B.6)
$=E_{\mu}^{\pi,\kappa}\sum_{a\in A}\sum_{s^{\prime}\in S}\Delta(a|X_{k-1})\kappa_{k-1}(s^{\prime}|H_{k-1},a)[r(X_{k-1},a)+\gamma u^{*}(s^{\prime})]$
$=E_{\mu}^{\pi,\kappa}\sum_{a\in A}\sum_{s^{\prime}\in S}\Delta(a|X_{k-1})K_{k-1}(s^{\prime}|X_{k-1},a)[r(X_{k-1},a)+\gamma u^{*}(s^{\prime})]$
$=E_{\mu}^{\pi,\kappa}\left(E_{\Delta(\cdot|s),K_{{k-1}(\cdot|s,\cdot)}}[r(s,A_{0})+\gamma u^{*}(X_{1})]\right)_{s=X_{k-1}}.$
where the last equality is a just a change of notation. Note that as $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$, we have $K_{k-1}(\cdot|X_{k-1},\cdot)\in\mathcal{P}_{X_{k-1}}$ w.p.1. Therefore,
$E_{\mu}^{\pi,\kappa}\left[r(X_{k-1},A_{k-1})+\gamma u^{*}(X_{k})\right]$ $\geq E_{\mu}^{\pi,\kappa}\left(\inf_{p_{s}\in\mathcal{P}_{s}}E_{\Delta(\cdot|s),p_{s}}[r(s,A_{0})+\gamma u^{*}(X_{1})]\right)_{s=X_{k-1}}$ (B.7)
$\stackrel{{\scriptstyle(i)}}{{\geq}}E_{\mu}^{\pi,\kappa}u^{*}(X_{k-1})-\delta$
where $(i)$ follows from (B.4).
Fix any $\epsilon>0$, let $\pi$ be defined as above. Recall the definition of $v_{\mu}^{\pi}(T,\kappa)$ in (A.2). We still choose $T$ large so that for all $\kappa$, $|v_{\mu}^{\pi}(T,\kappa)-v(\mu,\pi,\kappa)|\leq\epsilon/2$. Then, for any $\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$,
$\stackrel{{\scriptstyle(i)}}{{\geq}}E_{\mu}^{\pi,\kappa}\left[\sum_{k=0}^{T-1}\gamma^{k}r(X_{k},A_{k})\right]+\gamma^{T-1}\left[E_{\mu}^{\pi,\kappa}u^{*}(X_{T-1})-E_{\mu}^{\pi,\kappa}r(X_{T-1},A_{T-1})-\delta\right]$
$=v_{\mu}^{\pi}(T-1,\kappa)-\gamma^{T-1}\delta$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}\dots$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}v_{\mu}^{\pi}(0,\kappa)-\sum_{k=0}^{T-1}\gamma^{k}\delta$
$\geq E_{\mu}[u^{*}(X_{0})]-\frac{\delta}{1-\gamma},$
where $(i)$ applies inequality (B.7) and $(ii)$ iteratively repeats the previous steps.
By definition of $\inf$ and that $\pi\in\Pi_{\mathrm{S}}$, there exists $\kappa^{\prime}\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}$ s.t.
$v(\mu,\pi,\kappa^{\prime})-\frac{\epsilon}{4}\leq\inf_{\kappa\in\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}}v(\mu,\pi,\kappa)\leq v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}}).$ (B.8)
Then,
$0$ $\stackrel{{\scriptstyle(i)}}{{\leq}}E_{\mu}[u^{*}(X_{0})]-v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ (B.9)
$\stackrel{{\scriptstyle(ii)}}{{\leq}}E_{\mu}[u^{*}(X_{0})]-v(\mu,\pi,\kappa^{\prime})+\frac{\epsilon}{4}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}E_{\mu}[u^{*}(X_{0})]-v^{\pi}_{\mu}(T,\kappa^{\prime})+\frac{3\epsilon}{4}$
$\leq\frac{\delta}{1-\gamma}+\frac{3\epsilon}{4}.$
where $(i)$ follows from Proposition 2, $(ii)$ uses inequality (B.8), and $(iii)$ is due to the choice of $T$. Since $\delta$ can be chosen as $(1-\gamma)\epsilon/4$, this implies $E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})$ as $\epsilon>0$ is arbitrary.
B.2 Proof of Corollary 1.1
Proof.
Proof The uniqueness follows from the same contraction argument as in the proof of Proposition 1. Moreover,
$u^{*}(s)=\sup_{\phi\in\mathcal{Q}^{\mathrm{D}}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}\left[r(s,A_{0})+\gamma u^{*}(X_{1})\right]=\max_{a\in A}q^{*}(s,a).$
Plug the above equality into (3.3), one sees that $q^{*}$ is indeed the solution, hence equivalently solves (3.4). Lastly, Theorem 1 still holds in the context of $\mathcal{Q}=\mathcal{Q}^{\mathrm{D}}$. This implies the statement of the corollary. ∎
B.3 Proof of Theorem 2
We recall that in the setting of Theorem 2, $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$ can be an arbitrary subset. We highlight a consequence of the extra assumption (3.2) on $u^{*}$, namely the largest maxmin control value $v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$ is upper bounded by $E_{\mu}[u^{*}(X_{0})]$. Recall that by Proposition 3, we always have that the smallest maxmin control value $v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})=E_{\mu}[u^{*}(X_{0})]$. Together, these facts would imply the first item in Theorem 2.
@@ B.3 Proof of Theorem 2
Assume that the solution $u^{*}$ to (3.1) further satisfies (3.2), then $v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\leq E_{\mu}[u^{*}(X_{0})]$.
The proof to this proposition is deferred to Appendix B.3.1. Assuming the validity of this proposition, we proof of item 1 in Theorem 2.
Proof.
Proof of Item 1 in Theorem 2 Note that, by Lemma 1,
$v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v(\mu,\Pi,\mathrm{K})\leq v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})$
for all $\mu\in\mathcal{P}(\mathcal{S})$ and any mixture of $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\Pi=\Pi_{\mathrm{H}},\Pi_{\mathrm{M}},\Pi_{\mathrm{S}}$. This, Theorem 1, and Proposition 4 implies that
$E_{\mu}[u^{*}(X_{0})]=v(\mu,\Pi_{\mathrm{S}},\mathrm{K}_{\mathrm{H}}^{\mathrm{S}})\leq v(\mu,\Pi,\mathrm{K})\leq v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}})\leq v(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}).$
@@
Next we prove item 2 of Theorem 2. We will make use of the following proposition:
@@
$\sup_{\pi\in\Pi_{\mathrm{H}}}v(\mu,\pi,\kappa)\leq E_{\mu}[u^{*}(X_{0})]+\epsilon.$
The proof of Proposition 5 is deferred to the appendix Section B.3.1 as well. Equipped with Proposition 5, we can establish item 2 of Theorem 2.
@@
Proof of Item 2 in Theorem 2 Note that by assumption, weak duality of the sup-inf interchange and set inclusions, it is clear that
$E_{\mu}[u^{*}(X_{0})]=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}v(\mu,\pi,\kappa)\leq\inf_{\kappa\in\mathrm{K}}\sup_{\pi\in\Pi}v(\mu,\pi,\kappa)\leq\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}}\sup_{\pi\in\Pi_{\mathrm{H}}}v(\mu,\pi,\kappa)$ (B.10)
for any mixture of $\mathrm{K}=\mathrm{K}_{\mathrm{H}}^{\mathrm{S}},\mathrm{K}_{\mathrm{M}}^{\mathrm{S}},\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}$ and $\Pi=\Pi_{\mathrm{H}},\Pi_{\mathrm{M}},\Pi_{\mathrm{S}}$. On the other hand, the statement of Proposition 5 implies that
$\inf_{\kappa\in\mathrm{K}_{\mathrm{S}}^{\mathrm{S}}}\sup_{\pi\in\Pi_{\mathrm{H}}}v(\mu,\pi,\kappa)\leq E_{\mu}[u^{*}(X_{0})].$
Therefore, we conclude that the inequalities in (B.10) must be equalities. This proves item 2 of Theorem 2. ∎
B.3.1 Proof of Propositions 4 and 5